OpenStudy (anonymous):

-8x-5y=-1 find the value of y when x equals 2

OpenStudy (anonymous):

Put in 2 for x -8(2)-5y=-1 -16-5y=-1 -5y=15 y=-3

OpenStudy (anonymous):

thx @zeinas.

OpenStudy (anonymous):

You're welcome